OpenStudy (anonymous):

What is an expression that shows 11+7=5+13

OpenStudy (acacia):

Do you have multiple choice?

OpenStudy (anonymous):

Yes

OpenStudy (acacia):

Is one of them (5+6)+7=5+(6+7)? If so its that one because that's what I got 5 + 6 = 11 6 + 7 = 13 11+7 = 5 + 13

OpenStudy (anonymous):

Yes it is thank you!

OpenStudy (acacia):

You're welcome (: